Write the expected major organic product for 1-propynyl lithium (shown) with each of the following molecules. HyC-CEC:L: (aCHPlease be as detailed as possible!
Answer #1

when acetylide ion reacts with alkyl halide and forms the SN2 product as follows

HC-CECOL® + CH3CH2=CH-I → HC-CEC-CHEMACH, - Acetylece kan PAlkyl Holice a C–C=C tº + C CH=CH2 - H2C-CEC-CH (hoz CH₂ - 2° Alla